TWO PLANETS A AND B HAVE SAME MATERIAL DENSITY.IF RADIUS OF A IS TWICE THAT OF B, THEN RATIO OF ESCAPE VELOCITIES OF A AND B IS ?

Grade:11

1 Answers

Ramesh V
14 years ago

Ve = sqrt (2GM/r)

V2*r = constant

Va2*ra= Vb2*rb

2*Va2= Vb2

Va/Vb = 2-1/2

--

regards

Ramesh